### v1

# Performance evaluation of compact plastic scintillating fiber modules for muon tomography applications

Abstract
Muon tomography is a non-destructive imaging technique that exploits cosmic-ray muons from multiple directions. Its performance critically relies on stable, large-area, and high-resolution position-sensitive detectors. In this work, we report on the development of four compact scintillating fiber modules, each 100 cm long and composed of two staggered layers of 1 mm diameter fibers. The fibres are read out at one end by one-dimensional silicon photomultiplier arrays with a 2 mm pitch, coupled to Citiroc1A-based front-end electronics. The modules were characterised with cosmic-ray muons, yielding a detection efficiency above 97% and a mean spatial resolution of about 0.56 mm, with uniform response over different distances from the readout end. These results demonstrate the suitability of this detector design for compact and large-area systems in muon tomography applications.
1 Introduction
Cosmic-ray muons are secondary particles naturally produced by the interactions of primary cosmic rays with the Earth’s atmosphere. At sea level, their flux is typically about $\sim$$\rm{1\,cm^{-2}\,min^{-1}}$, with an average energy of $\sim$4 GeV [1]. Owing to their strong penetrating power, muon tomography can non-destructively reconstruct the internal structures of large and dense objects by tracking these muons from multiple directions [2, 3, 4]. Depending on the physical principle, the technique can be classified into muon absorption tomography, which exploits the muon transmission rate, and muon scattering tomography, which relies on the measurement of muon deflection angles. In recent years, muon tomography has been successfully applied in diverse fields, including archaeology [5], geological exploration [6], nuclear waste monitoring [7], and border security [8].
Given the relatively low flux of cosmic-ray muons and the complexity of practical application scenarios, muon tomography systems demand detectors that combine high spatial resolution with large-area coverage and long-term operational stability. Plastic scintillator detectors are often preferred in such applications, as they can be flexibly manufactured in various geometries, maintain stable performance across different temperatures, offer high detection efficiency, and allow relatively low-cost large-area production. To achieve good spatial resolution, scintillators are usually cut into long bars with rectangular [9] or triangular [10, 11] cross sections and read out with wavelength-shifting fibers. New 3D printing techniques have also been developed to manufacture scintillators with more complex geometries [12]. However, the spatial resolution of such scintillator detectors is typically at the millimeter level, and further improvement toward the sub-millimeter scale is difficult due to the constraints imposed by the bar shape and lateral size.
The plastic scintillating fiber (SciFi) detector offers much finer granularity compared to bulk scintillator detectors, enabling superior spatial resolution. In addition, its long attenuation length and high light yield make it possible to construct SciFi detectors with lengths of several meters. These features render the SciFi detector a promising candidate for compact muon tomography systems. This potential has been demonstrated in Refs. [13, 14] with round scintillating fibers, where grouped fibers were coupled with MAPMT or SiPM arrays to reduce the number of readout channels. To further enhance compactness and eliminate the laborious fiber-grouping process, we propose a SciFi detector employing a one-dimensional SiPM array readout [15] combined with a potential SiPM multiplexing strategy [16] for sub-millimeter spatial resolution. In this work, we focus on the design and performance evaluation of SciFi modules with SiPM array readout, while the multiplexing strategy will be discussed in a separate study.
The remainder of this paper is organized as follows. Section 2 describes the design of the 100-cm-long SciFi modules aiming for sub-millimeter spatial resolution. Section 3 presents the experimental setup and data acquisition system used for performance evaluation. Section 4 details the data analysis procedures, including pedestal substraction, gain calibration, signal clustering, tracking, and alignment. Section 5 reports the experimental results on the detection efficiency and spatial resolution of the SciFi modules at different positions. Finally, Section 6 concludes the work and provides an outlook for future developments.
2 Construction of SciFi modules
Figure 1: The design of one SciFi module with two staggered layers of 1mm diameter fibers and one-dimensional SiPM array of 2 mm pitch size for sub-millimeter spatial resolution.
The spatial resolution of SciFi detectors with one-dimensional SiPM array readout is generally determined by both the fibre diameter and the SiPM pitch. To reach sub-millimetre resolution, the preferred configuration is 1 mm diameter fibres coupled to SiPM arrays with a $1\times 2$ mm or $1\times 3$ mm size. However, due to the unavailability of such commercial SiPM products, the present SciFi modules are designed as shown in Fig. 1, employing two staggered layers of 1 mm diameter fibres read out by SiPM arrays with a $2\times 2$ mm size.
Figure 2: (a) Aluminium alloy mold with machined grooves, (b) first fiber layer placed along grooves of the mold, (c) the complete fiber modules with one end directly coupled to SiPM arrays and the other end covered by ESR film to reflect photons to the readout end, (d) the cross section of a SciFi module with two staggered layers of 1mm diameter fibers, and (e) one-dimensional SiPM array composed of 22 Hamamatsu S13360-2050VE devices used at the readout end.
Each compact SciFi module has dimensions of $5.5\times 100$ cm2 and is constructed using double-clad Kuraray SCSF-78M fibres [17]. These fibers have a typical trapping efficiency of about 5.4%, an attenuation length exceeding 4 m, and a light yield of roughly 8000 photons/MeV. Their emission spectrum spans 400–700 nm, peaking near 450 nm. To ensure precise positioning along the full module length, an aluminium alloy mould with precision-machined grooves is fabricated (Fig. 2(a)). After applying a release agent, the first fibre layer is placed in the grooves (Fig. 2(b)), followed by an adhesive mixed with $\rm TiO_{2}$ powder to reduce cross talks between fibers. The second layer is then positioned in the interstitial grooves. The assembly is covered with Teflon sheets, and demoulded once the adhesive solidifies. The fibre ends are cut and polished to ensure efficient optical coupling. Each module is mounted on a carbon-fibre support plate, with the readout end directly coupled to the SiPM arrays and the opposite end covered with enhanced specular reflector (ESR) film via an air gap (Fig. 2(c)). In addition, four SciFi modules of 10 cm length are produced using the same procedure but without ESR film to provide muon triggers in later experimental test.
To ensure efficient photon collection from the fibres, the dead area between adjacent SiPMs must be kept below 1 mm and minimized as much as possible. For this reason, Hamamatsu S13360-2050VE devices [18] with a $\rm 2\times 2$ mm2 photosensitive area were selected to assemble the one-dimensional arrays as shown in Fig. 2 (e). These devices employ through-silicon via (TSV) technology instead of conventional wire bonding, reducing the distance between the package edge and the photosensitive area to 0.2 mm on all four sides. An additional 0.1 mm margin was reserved to facilitate precise alignment and reliable soldering. The SiPMs are sensitive over the 320–900 nm wavelength range, matching well with the emission spectrum of the scintillating fibres. Their typical breakdown voltage is about 53 V and have peak photon detection efficiency reaches about 40% at 3 V overvoltage around 450 nm. Furthermore, each channel has a 50 $\rm\mu m$ pixel pitch and 1584 pixels, ensuring a linear response to photons emitted by the fibres when traversed by cosmic-ray muons.
3 Experimental setup
Figure 3: Experimental setup of the SciFi modules with Citiroc1A-based readout electronics. The four short modules provide the muon trigger for the long SciFi modules during cosmic-ray muon test. Their positions can be adjusted to record muon events at different locations along the long modules.
In the experimental setup shown in Fig. 3, four long SciFi modules of 100 cm length (layer 1-4), are arranged vertically with a spacing of approximately 5 cm along the $z$-axis and sensitive to $y$ positions of particles. Four shorter modules (layers 5–8), each 10 cm in length, are positioned perpendicularly to the long modules. Each short module can be slid along the $x$-axis to select muon trajectories at different positions along the long modules. The closest pair of short and long modules is separated by a vertical gap of about 0.5 mm and used to measure the two-dimensional position of muons. The long and short modules are read out by separate DT5550W data acquisition boards from CAEN, each equipped with four Citiroc 1A chips[19] on a piggyboard and a 14-bit ADC. The SiPM array of each SciFi module is connected to a dedicated Citiroc 1A chip and biased at 57 V via the piggyboard’s power supply module.
The Citiroc 1A is a 32-channel front-end ASIC optimized for SiPM readout. Each channel includes two parallel preamplifiers, one in high-gain (HG) mode and the other in low-gain (LG) mode, extending the dynamic range up to about 2500 photoelectrons (p.e.). Each preamplifier output is processed by a slow shaper with an adjustable shaping time between 12.5 ns and 87.5 ns, followed by a peak-sensing circuit to measure the integrated charge. For triggering, the chip provides a 15ns peaking-time fast shaper capable of discriminating signals down to 1/3 p.e. in each channel, as well as an ASIC-level charge trigger formed by the logical OR of all 32 channels. The DT5550W board can then generate a global trigger by combining the logic OR outputs of its four Citiroc 1A chips.
During the experimental tests of the SciFi modules, the HG signal was amplified by a factor of 26 and the LG signal by a factor of 2.6, with the slow shaping time of all channels fixed at 87.5 ns. Three dedicated data acquisition modes were implemented for pedestal estimation, gain calibration, and cosmic-ray muon detection, respectively:
  • Pedestal mode: SiPM signals are recorded without bias voltage, using 1 kHz periodic triggers generated by the DT5550W software.
  • Gain-calibration mode: The two DT5550W boards operates independently. A global trigger is issued whenever the charge in any channel of the four Citiroc 1A chips exceeds about 5 p.e., in which case all SiPM signals from either the four long or the four short modules were read out.
  • Muon-test mode: The DT5550W board for the short modules operates with ASIC-level charge trigger, issuing a global trigger whenever a channel of any Citiroc1A chip exceeded about 10 p.e. The threshold is chosen as a compromise between suppressing background events and retaining muon events. This trigger is sent to the DT5550W board of the long modules as an external trigger. Its trigger latency, including cables, is measured to be about 40 ns using charge injection into a Citiroc 1A under identical acquisition settings. This ensures that the trigger arrives before the peak-sensing stage of the long modules with 87.5 ns shaping time. Upon receiving the external trigger, the long-module data are read out. Complete events are finally built by matching long- and short-module data with the same trigger ID in offline analysis.
4 Data analysis
4.1 Channel calibration
Although in the cosmic-ray muon test all SiPM channels of a SciFi module were biased at the same voltage, variations in their breakdown voltages and the responses of different Citiroc1A channels resulted in non-uniform responses of number of photonelectrons ($N_{\mathrm{pe}}$) among SiPM channels. To account for these variations, the pedestal must be subtracted from the raw data and the gain of each SiPM channel calibrated individually.
Refer to caption Refer to caption
Figure 4: (Left) The raw ADC spectrum of one HG amplifier channel to extract gain and pedestal values. (Right) LG signal after pedestal substraction as a function of the calibrated HG $N_{\mathrm{pe}}$ for a representative SiPM channel, from which the LG gain is obtained by the linear fit.
As shown in the left panel of Fig. 4, the raw ADC spectrum of the HG amplifier signals obtained in gain-calibration mode exhibits a clear offset from zero. The pedestal is defined as the ADC value of the first and highest peak (from left to right), which is consistent with the value measured in pedestal mode with no high voltage applied to the SiPM arrays. The same procedure is applied for the pedestal estimation of the LG amplifier signals. The charge threshold set during gain-calibration mode corresponds to the valley at approximately 5 p.e. The peaks between the pedestal and the threshold are identified, and the HG gain, $G_{HG}$, is determined as the mean ADC separation between adjacent peaks. As reported in Refs. [20, 21], the ADC difference between the pedestal and the 1 p.e. peak is systematically smaller than the actual gain, since the peak-sensing stage always samples the maximum signal value within the hold-delay window. Consequently, the 1 p.e. peak can be treated as a “fake pedestal” to estimate the number of photoelectrons in an HG signal as
$N_{\mathrm{pe}}=\frac{ADC_{HG}-ADC_{1\mathrm{pe}}}{G_{HG}}+1,.$ (4.1)
The LG gain is intercalibrated with respect to the HG signals, as illustrated in the right panel of Fig. 4. It is obtained as the slope of a linear fit using HG signals in the range of 10–90 p.e. The ratio of obtained HG gain and LG gain as shown in Fig. 4 is consistent with that of amplifier value set in firmware. For larger SiPM signals (>90>90 p.e.), the HG response becomes non-linear, and in such cases the LG signals are used to estimate the input $\rm{N_{pe}}$. After this calibration, the raw ADC data acquired in muon-test mode are converted into $N_{\mathrm{pe}}$ values. As shown in the left panel of Fig. 5, all 22 SiPM channels of layer 1 exhibit consistent signal spectra after pedestal subtraction and gain calibration. Moreover, the distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event is consistent across all four long modules, as illustrated in the right panel of Fig. 5. The slight deviation observed in layer 2 is attributed to differences in the procedure used when attaching the ESR films.
Refer to caption Refer to caption
Figure 5: (Left) Calibrated SiPM signals of layer 1 acquired in muon-test mode after pedestal subtraction and gain calibration. (Right) Distribution of the SiPM channel with the maximum $N_{\mathrm{pe}}$ in each event for all four long modules when the short modules were centered at 52.5 cm from the readout end of the long modules.
4.2 Signal clustering and track reconstruction
Cosmic-ray muons passing through a SciFi module may deposit energy in multiple fibers, thereby inducing SiPM signals in one or more neighboring channels. A group of such neighboring channels is defined as a cluster. To select muon signals while suppressing background events from ambient radiation, the following procedure is applied: the seeding channel is identified as the one with a signal exceeding 8 p.e., corresponding to the valley in the $N_{\mathrm{pe}}$ distribution shown in the right panel of Fig. 5. Neighboring channels with signals above 6 p.e. are then included together with the seeding channel to form a cluster. The number of SiPM channels contained in a cluster is referred to as the cluster size, clscls. The hit position of the cluster on the SciFi module is determined using the center-of-gravity (CoG) method:
$x_{\mathrm{cog}}=\frac{\sum_{i}N_{i}x_{i}}{\sum_{i}N_{i}},$ (4.2)
where $x_{i}$ is the center position of the $i$th SiPM in the cluster, and $N_{i}$ is its corresponding signal.
To suppress background clusters or clusters contaminated by background channels, only those with a cluster size $cls\leq 3$ in one layer are retained for track reconstruction. Particle tracks are reconstructed separately in the $xz$ and $yz$ planes. All possible combinations of clusters across different layers are scanned and tested with a straight-line fit by the least square method, and the set of clusters yielding the smallest fitting $\chi^{2}$ is selected as the track candidate for that event. Fig. 6 illustrates an event display of reconstructed tracks in the $xz$ and $yz$ projection planes after alignment as explained in Sec. 4.3.
Figure 6: Event display of reconstructed tracks in the $xz$ and $yz$ projection planes. The blue rectangles represent SiPM arrays, while the red rectangles indicate the signal cluster with black numbers as their $N_{\mathrm{pe}}$ values in each channel.
4.3 Module alignment
The fitting residual of a SciFi module, defined as the difference between the reconstructed cluster position and the extrapolated position from the fitted track onto the same layer, originates from the intrinsic detector resolution, multiple-scattering effects, and mechanical tolerances of the setup assembly. To obtain a more precise determination of the intrinsic position resolution, as discussed in Sec. 5.2, a track-based alignment procedure is applied to all modules in the experimental setup.
Figure 7: The residual distributions of all SciFi modules before and after alignment.
In this procedure, a transverse translation $s$ and a small rotation angle $\theta$ around the $z$-axis are considered as free alignment parameters for each SciFi module. The alignment constants are determined by minimizing the global sum of squared residuals,
$R^{2}=\sum_{N}\sum_{i}\left[\left(y_{i}^{c}-y_{i}^{f}\right)^{2}+\left(x_{i}^{c}-x_{i}^{f}\right)^{2}\right],$ (4.3)
where $N$ is the number of tracks, $x_{i}^{c}$ and $y_{i}^{c}$ are the corrected cluster positions of the nearest pair of short and long modules after applying the alignment parameters, and $x_{i}^{f}$ and $y_{i}^{f}$ are the extrapolated positions of the tracks fitted by corrected cluster positions in the $xz$ and $yz$ planes, respectively. The corrections are expressed as
$y_{i}^{c}$ $=y_{i}^{m}+x_{i}^{m}\cdot\theta_{y_{i}}+s_{y_{i}},$ (4.4)
$x_{i}^{c}$ $=x_{i}^{m}-y_{i}^{m}\cdot\theta_{x_{i}}+s_{x_{i}},$
where $x_{i}^{m}$ and $y_{i}^{m}$ denote the measured cluster positions in the paired short and long modules, respectively.
During the alignment procedure, only the cluster with the maximum signal on each layer is used in track fitting. Tracks are required to satisfy the following criteria: a cluster size of $cls\leq 3$ on all four layers, and a fit quality of $\chi^{2}<100$. The tracks reconstructed with cluster of $cls=1$ on all four long or short modules are dimissed to avoid biased residuals as explained in Sec. 5.2. The minimization of Eq. 4.3 is performed iteratively with the Gauss–Newton method. For each test position of SciFi modules, at least 1000 selected tracks are used for the alignment. Covergence to level of 0.1 mm of translation parameters is realized typically after 3 iterations. Fig. 7 shows an example of the residual distributions with tracks of $\chi^{2}<10$ of all SciFi modules before and after alignment using muon test data, clearly demonstrating that the mean of the residuals shifts to zero after alignment.
5 Results
The detection efficiency and spatial resolution are key performance parameters of the SciFi modules, as they determine both the imaging time and the accuracy of muon tomography. These parameters are evaluated using muon events recorded at five different distances between the center of the short module and the readout end of the long SciFi modules along the $x$ axis. The different measurement positions were achieved by sliding the short modules within the experimental setup. For each position, data are collected over periods longer than five days in muon-test mode.
5.1 Detection efficiency
To evaluate the detection efficiency of a long SciFi module under test (DUT) at a given measurement position, good events are selected according to the following criteria:
  • The short modules, excluding the one paired with the DUT, are required to reconstruct $xz$-plane tracks with $\chi^{2}<2$ and cluster sizes of $cls\leq 3$ in each participating layer. This ensures that the selected muons traversed the intended measurement position on the long modules.
  • The long modules, excluding the DUT, are required to reconstruct $yz$-plane tracks with $\chi^{2}<2$ and cluster sizes of $cls\leq 3$ in each participating layer. In addition, the extrapolated track position on the DUT is required to lie within its sensitive area.
Figure 8: Example distribution of cluster size versus cluster signal $N_{\mathrm{pe}}$ for good events in layer 2, measured at a position of 52.5 cm.
With these criteria, about 1000 good events per day are obtained for each DUT. Fig. 8 shows, as an example, the cluster size and cluster signal distributions for the DUT in layer 2 at a measurement position of 52.5 cm. Most clusters have $cls\leq 2$, and the most probable value of $N_{\mathrm{pe}}$ is about 36 p.e., consistent with the module design expectation. Clusters with larger sizes are typically associated with higher signals, which can originate from muon hit combined with secondary particles produced in muon interactions, such as $\delta$ rays, or with ambient radiation.
For each good event, the DUT is considered to have detected the muon if a cluster with $cls\leq 3$ is present on the DUT layer and the residual between the cluster position and the extrapolated track position in the $yz$ plane is less than 5 mm. Events with $cls=3$ are retained to maximize statistics while only marginally affecting the position resolution. The residual threshold is particularly relevant for the outermost modules, where it is influenced by the geometrical layout. The detection efficiency is then defined as
$\epsilon=\frac{N_{\mathrm{det}}}{N_{\mathrm{good}}}\;,$ (5.1)
where $N_{\mathrm{det}}$ is the number of events with a valid cluster on the DUT and $N_{\mathrm{good}}$ is the total number of good events. Fig. 9 shows the measured detection efficiencies of the four long SciFi modules at different measurement positions. The efficiencies are uniform along the module length and consistently exceed 97%.
Figure 9: Detection efficiency of the four long SciFi modules measured at different positions along their length. Error bars represent statistical uncertainties.
5.2 Spatial resolution
The spatial resolution of a DUT is evaluated from the residual distribution between the detected cluster position on the DUT and the extrapolated track reconstructed with the other layers. The measured residual variance, $\sigma_{r}$, includes contributions from both the intrinsic spatial resolution of the DUT, $\sigma_{d}$, and the resolution of the extrapolated track, $\sigma_{f}$:
$\sigma_{r}^{2}=\sigma_{d}^{2}+\sigma_{f}^{2}\;.$ (5.2)
Since tracks are reconstructed with a straight-line fit, the extrapolated hit position is a linear combination of the reference layers. Accordingly, the resolution of the extrapolated position at module $m$ can be expressed as [22]:
$\sigma_{f,m}^{2}$ $=\sum_{j\neq m}w_{mj}\sigma_{d,j}^{2},$ (5.3)
$w_{mj}$ $=\frac{(z_{j}-\bar{z})(z_{m}-\bar{z})}{\sum_{j\neq m}(z_{j}-\bar{z})^{2}}+\frac{1}{N-1},$
$\bar{z}$ $=\frac{1}{N-1}\sum_{j\neq m}z_{j}\;,$
where $N$ is the total number of long modules and $z_{j}$ denotes the vertical position of module $j$ along the $z$ axis. The spatial resolution of each module was then obtained by solving the set of linear equations for all four modules.
For the detected good events defined in Sec. 5.1, the residual variance is usually determined by fitting the residual distribution with a Gaussian function. However, as shown in Fig. 10(a), the residual distribution of the DUT in layer 2 exhibits a Gaussian-like shape for clusters with $cls>1$, whereas discrete spikes are observed for clusters with $cls=1$.
[…]
Figure 10: Residual distributions of the DUT in layer 2: (a) for all detected good events with clusters of $cls=1$ or $cls>1$ on the DUT, (b) for good events with extrapolated tracks reconstructed exclusively from $cls=1$ clusters in the three reference layers, and (c) for good events after excluding tracks reconstructed exclusively from $cls=1$ clusters in the three reference layers, together with a Gaussian fit of their residual distribution.
Because the SiPM pitch is larger than the fiber diameter, about 68% of muon clusters correspond to $cls=1$, as shown in Fig. 8. In these cases, the reconstructed cluster positions by the CoG method are fixed at the SiPM channel centers. As a result, roughly 30% of tracks are reconstructed exclusively from $cls=1$ clusters in only three reference layers. Such tracks with hits on three SiPM centers lead to systematic reconstruction biases imposed by the fixed detector layout of SiPMs as illustrated in Fig. 6. Furthermore, since only DUT clusters with residuals smaller than 5 mm are accepted, the residual distribution of such biased tracks shows discrete spikes for $cls=1$ clusters, as seen in Fig. 10(b). Similar patterns are also observed for other layers.
After excluding tracks reconstructed solely from $cls=1$ clusters in the reference layers, the combined residual distribution of DUT clusters with $cls=1$ and $cls>1$ is well described by a Gaussian fit, from which $\sigma_{r}$ is extracted (Fig. 10(c)). Although the intrinsic resolution is expected to be better for $cls>1$ clusters than for $cls=1$, in analogy with the performance of segmented pixel detectors in Ref. [23], here we report only the overall spatial resolution constrained by the limited number of experimental layers and muon events. Fig. 11 shows the extracted spatial resolutions of all four SciFi modules at different measurement positions. The average resolution across all layers and positions is about $0.56\pm 0.04$ mm. The modules exhibit an approximately uniform response along their length.
Figure 11: Spatial resolution of the four 100-cm long SciFi modules measured with muons at different distances from the readout end. The error bars are obtained from the propagation of fitting uncertainties through the solution of the linear equations. The shaded band indicates the average spatial resolution across all layers and measurement positions.
6 Conclusion
In this work, we designed and constructed four compact 100 cm long SciFi modules, each comprising two staggered layers of 1 mm diameter plastic scintillating fibers read out at one end by a one-dimensional SiPM array with a 2 mm pitch. The modules were characterized using cosmic-ray muons triggered by 10 cm short SciFi modules positioned at different distances along their length, with data acquired by Citiroc 1A–based electronics. After dedicated channel calibration, signal clustering, track reconstruction, and alignment, the modules achieved detection efficiencies above 97% and an average spatial resolution of about 0.56 mm. Both parameters were found to be uniform across the full module length.
Compared with previously developed scintillating fiber detectors for muon tomography, the present design is more compact and avoids the laborious fiber-grouping process, while providing sub-millimeter spatial resolution superior to that of bulk scintillator detectors. These features demonstrate the potential of the proposed modules for building compact, high-resolution muon tomography systems. Future work focusing on SiPM readout multiplexing strategies is under development to reduce the number of electronics channels, thereby advancing toward scalable and cost-effective large-scale applications.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 12205174.